OpenStudy (anonymous):

3. Write an equation of the line that passes through (6, -2) with slope -3. A. y = 6x - 3 B. y = -3x C. y = -2x + 6 D. y = -3x +16

OpenStudy (anonymous):

The equation of any line is of the form y=mx+b "m" is the slope and "b" is the initial value. Obviously, it's B, C or D since the slope is -3. So your equation is of the form y=-3x+b. Now you need to find "b". You have a point (x,y): (6,-2). You can plug that into your equation and slove for b. y=-3x+b (-2)=-3(6)+b

OpenStudy (anonymous):

d

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Yup!

Parth (parthkohli):

An easier way to use is the point-slope form:\[y - y_1 = m(x - x_1)\]So you have the point \((6,-2)\) and the slope \(-3\). You can write the point-slope in this case as follows:\[y - (-2) = -3(x - 6)\]Can you simplify that? :)

OpenStudy (anonymous):

thank you everyone!!!

OpenStudy (anonymous):

Parth, students don't usually that form of a line... I would try to stick to the method they learn not to confuse them :P

Parth (parthkohli):

@BluFoot: Haha, okay... but this form is really convenient :)